### v1

# Vision-Language Models Represent Darker-Skinned Black Individuals as More Homogeneous than Lighter-Skinned Black Individuals

Abstract
Vision-Language Models (VLMs) combine Large Language Model (LLM) capabilities with image processing, enabling tasks like image captioning and text-to-image generation. Yet concerns persist about their potential to amplify human-like biases, including skin tone bias. Skin tone bias, where darker-skinned individuals face more negative stereotyping than lighter-skinned individuals, is well-documented in the social sciences but remains under-explored in Artificial Intelligence (AI), particularly in VLMs. While well-documented in the social sciences, this bias remains under-explored in AI, particularly in VLMs. Using the GAN Face Database, we sampled computer-generated images of Black American men and women, controlling for skin tone variations while keeping other features constant. We then asked VLMs to write stories about these faces and compared the homogeneity of the generated stories. Stories generated by VLMs about darker-skinned Black individuals were more homogeneous than those about lighter-skinned individuals in three of four models, and Black women were consistently represented more homogeneously than Black men across all models. Interaction effects revealed a greater impact of skin tone on women in two VLMs, while the other two showed nonsignificant results, reflecting known stereotyping patterns. These findings underscore the propagation of biases from single-modality AI systems to multimodal models and highlight the need for further research to address intersectional biases in AI.
1 Introduction
Large Language Models (LLMs), such as GPT-4, have rapidly advanced the fields of natural language understanding and generation, enabling applications in areas like automated content creation and decision support. These models are trained on extensive collections of text, providing them with remarkable capabilities in a wide array of language-related tasks. Vision-Language Models (VLMs) represent a step further in this technological advancement, integrating LLM capabilities with image processing tasks from image captioning to text-to-image generation.
With advancements of Large- and Vision-Language Models, concerns about their potential to reproduce and amplify human biases have intensified. LLMs, for instance, often generate text aligned with group-based stereotypes (e.g., Abid et al., 2021; Lucy and Bamman, 2021). Recent studies have extended this inquiry to VLMs, revealing that these models produce stereotypical captions and answers for image inputs (Zhou et al., 2022; Zhao et al., 2021) and generating biased images, such as lighter-skinned men as software developers and darker-skinned women as housekeepers (e.g., Bianchi et al., 2023; Naik and Nushi, 2023; Sun et al., 2023; Sami et al., 2023).
1.1 Homogeneity Bias in AI Models
Another form of stereotyping documented in AI models is homogeneity bias—the tendency of AI systems to represent certain groups as more uniform than others. Rooted in social psychology, this phenomenon relates to the out-group homogeneity effect, where individuals perceive their own group (in-group) as more diverse than other groups (out-group) (Linville et al., 1989; Quattrone and Jones, 1980).
Lee et al. (2024) explored whether a state-of-the-art LLM reproduces similar biases, examining how group status affects the homogeneity of group representations. They found that ChatGPT generated more uniform texts for racial/ethnic minority groups and women compared to the racial/ethnic majority group in the US (i.e., White Americans) and men. They suggested that disproportionate representation and stereotypical portrayals of marginalized groups in training data may lead LLMs to replicate this pattern of human stereotyping. Similarly, Cheng et al. (2023) found that texts about marginalized groups often exaggerate defining characteristics, resulting in more caricature-like representations.
These findings align with broader research on stereotyping and erasure in Natural Language Processing (NLP) systems, which highlights minimal representation and stereotypical portrayals of marginalized groups, leading to erasure—the failure to adequately represent the diversity and richness of an identity (Dev et al., 2022). For example, prior work has shown that contextualized word embeddings failed to provide meaningful representations for non-binary gender pronouns in the embedding space (Dev et al., 2021). Biases like these can perpetuate societal inequalities by reinforcing misrepresentation and stereotypes about marginalized groups. Furthermore, as these models become pervasive in everyday life, they risk wrongly influencing user perceptions. Evidence suggests that AI biases can shape attitudes and decision-making (e.g., Fisher et al., 2024), making homogeneity bias in AI models concerning for its potential to reinforce skewed perceptions and erasure.
1.2 Skin Tone Bias
Skin tone bias refers to prejudiced attitudes or discriminatory behaviors directed toward individuals within a racial group based on their skin tone (Maddox, 2004; Maddox and Gray, 2002). Research in this area has predominantly focused on perceptions of Black individuals, showing that darker-skinned Black individuals are more likely to be associated with negative traits compared to lighter-skinned Black individuals (e.g., Berumen Martinez, 2021; Stepanova and Strube, 2018; Kahn and Davies, 2011; Lynn, 2002; Maddox and Gray, 2002).
This bias also has significant real-world consequences. For example, darker-skinned Black individuals are receive lower ratings and fewer job offers in hiring scenarios (Wade et al., 2004; Harrison and Thomas, 2009), achieve lower levels of educational attainment and income (Keith and Herring, 1991), experience greater racial discrimination (Klonoff and Landrine, 2000), and report higher levels of mental distress due to discrimination (Gleiberman et al., 1995). Overall, skin tone plays a critical role in shaping both perceptions and life outcomes for Black individuals in the United States.
1.3 This Work
While substantial research in the social sciences has examined the impact of skin tone on social perceptions and real-world outcomes, relatively few studies have addressed skin tone bias in Artificial Intelligence (AI) models, particularly in Vision-Language Models (VLMs). Seminal work by Buolamwini and Gebru (2018) revealed that commercial gender-classification systems perform significantly better for lighter-skinned individuals. While much subsequent research has investigated the impact of skin tone on machine learning model performance (e.g., Groh et al., 2024; Kinyanjui et al., 2019), the topic of skin tone bias in newer generative models (e.g., VLMs) remains relatively under-explored.
Figure 1: Summary of the experimental setup. We collect 50-word stories about Black individuals differing in skin tone using four state-of-the-art Vision-Language Models. After encoding these stories into sentence embeddings, we compare the pairwise similarity of the embeddings using mixed-effects models.
Skin tone bias in Vision-Language Models (VLMs) carries significant implications. As VLMs increasingly contribute to the production of social information and creative content, they may generate more stereotypical texts in response to images of darker-skinned individuals compared to lighter-skinned individuals. Research shows that repeated exposure to stereotypes in social information can reinforce societal biases (Park et al., 2007; Saleem et al., 2017), suggesting that if VLMs perpetuate skin tone bias, they could exacerbate the very inequities that led to the bias in their training data. Moreover, as VLMs and other generative models are increasingly deployed in decision-making contexts, such biases could amplify discrimination in real-world outcomes.
We explored the topic of skin tone bias in VLMs through the lens of homogeneity bias, investigating how skin tone is associated with homogeneity of group representations. Our study was informed by hypotheses and research questions grounded in the social psychology literature on skin tone bias. Based on extensive evidence of the role of skin tone in stereotyping, we hypothesized that stories about darker-skinned Black individuals would exhibit greater homogeneity than those about lighter-skinned Black individuals. To test this, we encoded VLM-generated stories into sentence embeddings and measured their pairwise cosine similarity, following the method introduced by Lee et al. (2024). A mixed-effects model was then fitted with skin tone, derived from the images, as the only predictor. Additionally, we investigated the interaction between skin tone and gender as a research question. Prior studies suggest that skin tone effects may differ across genders, with stronger influences on attractiveness perceptions among women (Hill, 2002) and more pronounced performance disparities found in commercial gender-classification systems within women compared to men (Buolamwini and Gebru, 2018).
2 Method
2.1 Image Stimuli
We sampled ten images of Black American men and women from the publicly available GAN Face Database (GANFD) [^1] https://osf.io/7auyw, which features realistic, computer-generated faces. The database includes sets of images representing the same fictional individuals, with manipulations applied to vary facial features associated with perceived race. To ensure consistency, we used cropped images with a uniform grey background that contained only the face.
Figure 2: Two face stimulus pairs representing (A) Black men and (B) Black women. In each pair, the left image depicts a lighter-skinned individual, while the right shows a darker-skinned individual, generated from the same set.
From the GANFD, we focused on facial stimuli that were primarily identified as Black or multiracial. [^2] Multiracial faces were included due to an insufficient number of faces rated as Black. Within each set, faces with higher perceived Blackness ratings, measured by the question, “How much does this person look Black or African American? (0-100),” were assigned to the darker-skinned condition, while those with lower perceived Blackness scores were assigned to the lighter-skinned condition. As shown in Figure 2, faces with higher perceived Blackness typically had darker skin tones, while other facial features remained largely consistent across the skin tone conditions. This approach allowed us to manipulate skin tone while controlling for other features associated with perceived Blackness, such as nose width or eyebrow thickness (Ma et al., 2018).
2.2 Selection of Vision-Language Models and Writing Prompts
We collected data using a collection of Vision-Language Models including two proprietary VLMs—GPT-4o mini, GPT-4 Turbo—and two open-source VLMs—Llama-3.2-11B-Vision-Instruct (“Llama-3.2") (Grattafiori et al., 2024) and BLIP-3 (xgen-mm-phi3-mini-instruct-r-v1) (Xue et al., 2024). We accessed the proprietary models using the OpenAI API and the open-source models by downloading the model weights and running inferences on them locally.
The models were given the following writing prompt, “Write a 50-word story about the individual inside the image." and the following system prompt, “You are a helpful chat assistant. You are going to generate texts in response to images depicting fictional individuals." [^3] The system prompt was not supplied to Llama-3.2 as it did not support system-level instructions. The maximum number of generated tokens was set to 150, and 50 stories were generated for each image, totaling 1,000 completions from each VLM. All data collection involving open-source models was conducted using an NVIDIA RTX A6000 GPU.
2.3 Measure of Homogeneity
To quantify homogeneity of stories generated for each group, we adopted the measure introduced by Lee et al. (2024). We first represented the generated stories into sentence embedding representations using a pre-trained Sentence-BERT model (Reimers and Gurevych, 2019)—specifically all-mpnet-base-v2—and then calculated the cosine similarity between all possible combinations of sentence embeddings of stories generated for each image. Larger cosine similarity between sentence embeddings indicates that the stories are more similar to each other and, hence, more homogeneous.
2.4 Comparison of Cosine Similarity Measures
We fitted mixed-effects models (Bates et al., 2014; Pinheiro and Bates, 2000) to compare cosine similarity measurements between groups. These models account for random variations in measurements, such as those arising from repeated cosine similarity values for texts generated from images within the same condition (i.e., skin tone and gender), as well as variations in baseline similarity across conditions. To account for the resemblance between facial stimuli generated from the same set thereby possibly affecting the similarity of the generated stories, the sets of facial stimuli (i.e. Pair ID) were used as random intercepts in all our mixed-effects models.
First, we fitted a Skin Tone Model with skin tone as the sole fixed effect to test the hypothesis that stories about darker-skinned Black individuals are more homogeneous than those about lighter-skinned Black individuals. Next, we fitted a Gender Model with gender as the sole fixed effect to test the hypothesis that stories about women are more homogeneous than those about men. Finally, we fitted a Skin Tone and Gender Interaction Model with skin tone, gender, and their interactions to examine the interaction between skin tone and gender.
In all models, lighter skin tone was set as the reference level, with a significantly positive effect of skin tone indicating larger cosine similarity values for darker-skinned Black individuals than for lighter-skinned Black individuals. For gender, men were set as the reference level, with a significantly positive effect indicating higher cosine similarity values for Black women compared to Black men. Mixed-effects models were fitted using the lme4 package (Lenth et al., 2024), and skin tone effects within gender groups were evaluated using the emmeans package (Lenth et al., 2024). Likelihood-ratio tests, conducted with the afex package, assessed whether adding individual terms improved model fit. All analyses were performed in R version 4.4.0.
2.5 Power Analysis
We used 10 facial stimuli per group to assess homogeneity bias. Power analysis with the simr package in R (Green et al., 2023), employing Monte Carlo simulations, determined that 1,245 cosine similarity measurements per pair of stimuli (i.e., Pair ID) were required to achieve 90% power for detecting an interaction effect (skin tone $\times$ gender) with an effect size of 0.30 (from Lee et al. (2024)) at $\alpha$ = .05. With 2,500 cosine similarity measurements collected, our study exceeded this requirement. This analysis accounted for the increased power requirements of detecting interaction effects, ensuring that our study was adequately powered to test all effects with statistical confidence.
3 Results
In the Results section, we summarize the Skin Tone Model output to evaluate the effect of skin tone, presenting likelihood-ratio test results and a visualization of cosine similarity measurements for each skin tone group. We then summarize the Gender Model output to evaluate the effect of gender, presenting likelihood-ratio test results and a visualization of cosine similarity measurements for each gender group. Finally, we analyze the Interaction term from the Skin Tone and Gender Interaction Model, presenting likelihood-ratio test results, simple slopes, and a visualization of cosine similarity measurements for each intersectional group.
3.1 Darker Skin Tone Associated with Increased Homogeneity
Figure 3: Standardized cosine similarity values of lighter and darker skinned Black individuals generated from all four VLMs. Higher standardized cosine similarity means more homogeneity in the stories generated for that group. Error bars represent one standard error above and below the mean.
As shown in Figure 3, stories about darker-skinned Black individuals were significantly more homogeneous than those about lighter-skinned Black individuals in GPT-4o mini, GPT-4 Turbo, and Llama-3.2 (bs = 0.044, 0.15, and 0.080, respectively, ps < .001). However, we found an opposite pattern in BLIP-3 (bs = -0.095, ps < .001). Likelihood-ratio tests revealed that including skin tone improved model fit for all four VLMs ($\chi^{2}$(1)s $\geq$ 289.55, ps < .001). See Table S2 for summary output of the Skin Tone Models and Table S5 for likelihood-ratio test results.
3.2 VLMs Represent Women as More Homogeneous than Men
Figure 4: Standardized cosine similarity values of Black men and women generated from all four VLMs. Higher standardized cosine similarity means more homogeneity in the stories generated for that group. Error bars represent one standard error above and below the mean.
As shown in Figure 4, stories about Black women were significantly more homogeneous than those about Black men across all four models (bs = 0.63, 0.15, 0.40, and 0.24, respectively, ps < .001). Likelihood-ratio tests revealed that including gender improved model fit for all four VLMs ($\chi^{2}$(1)s $\geq$ 7.34, ps < .01). See Table S3 for summary output of the Skin Tone and Gender Interaction Models and Table S5 for likelihood-ratio test results.
3.3 Interaction between Skin Tone and Gender
Figure 5: Standardized cosine similarity values of lighter and darker skinned men and women generated from all four VLMs. Higher standardized cosine similarity means more homogeneity in the stories generated for that group. Error bars represent one standard error above and below the mean.
As shown in Figure 5, we found mixed evidence for the interaction between skin tone and gender. In GPT-4o Mini and Llama-3.2, a positive interaction effect indicated that the impact of skin tone on homogeneity was significantly greater for women than for men (bs = 0.10 and 0.21, ps < .001). However, in GPT-4 Turbo and BLIP-3, the interaction effect was not significant (bs = -0.0048 and -0.0072, ps = .038 and .028). Likelihood-ratio tests showed that including the interaction effect improved model fit for VLMs with significant positive interactions ($\chi^{2}$(1)s = 389.86 and 838.42, ps < .001) but not for those with non-significant interactions ($\chi^{2}$(1)s = 0.76 and 1.17, ps = .38 and .28). See Table S4 for summary output of the Skin Tone and Gender Interaction Models, Table S5 for likelihood-ratio test results, and Table S6 for simple slopes analysis results.
4 Discussion
4.1 Skin Tone Bias Persists in Vision-Language Models
With the vision modality of AI models, the question of whether skin tone of individuals given as input to these AI models influences their output poses a significant threat to both stereotyping and erasure. Despite past evidence that skin tone can affect performances of gender classification systems, such questions had not been explored with respect to the more recent AI models such as VLMs. In this work, we find evidence that skin tone affects homogeneity bias, a form of stereotyping found in LLMs, suggesting that VLMs process skin tone of individuals and associated semantic attributes based solely on skin tone.
Such biases related to skin tone may partly stem from the under-representation of darker-skinned individuals in training data, a documented source of performance disparities in other AI models (e.g., Buolamwini and Gebru, 2018). Given that VLMs are primarily trained on web-scraped data, which also suffers from similar under-representation issues (see Bender et al. (2021) for a detailed review), this attribution likely applies to homogeneity bias of VLMs as well.
4.2 Convergent Evidence of Gender Homogeneity Bias
Consistent with past evidence finding that women were represented as more homogeneous relative to men in an LLM (Lee et al., 2024), we found evidence of gender homogeneity bias in VLMs. This raises the possibility that bias found in AI models that have been trained on single modalities may be passed onto multimodal models, especially when the bias remains unattended. Future work would benefit from a systematic analysis of how different modalities interact to affect homogeneity bias, thereby clarifying if added modalities on top of text amplify biases and what parts of the model architecture would be most effective for targeting bias mitigation efforts in VLMs.
4.3 The Disproportionate Effect of Skin Tone on Women
In two of four VLMs—GPT-4 Turbo and Llama-3.2—the effect of skin tone on homogeneity of group representations was significantly greater for women than for men. Upon closer inspection of the Skin Tone and Gender Interaction Models, including the interaction effect to the Skin Tone models rendered the effect of skin tone either insignificant or in the opposite direction, suggesting that the main effect of skin tone in the Skin Tone models were primarily driven by the effect of skin tone within Black women. This, to some extent, is consistent with human stereotyping patterns identified by Hill (2002), where skin tone disproportionately affects women. While Buolamwini and Gebru (2018) demonstrated intersectional bias in gender classification systems, our results demonstrate that similar biases persist in Vision-Language Models (VLMs), reinforcing the need to address intersectional bias across AI systems.
5 Limitations
A limitation of this work is the potential conflation of perceived Blackness (or racial prototypicality of Black Americans) and skin tone. Face stimuli were assigned to lighter- and darker-skinned conditions based on perceived Blackness ratings from GANFD. Although we justified this approach by emphasizing that the manipulations of face stimuli primarily reflected changes in skin tone, it remains unclear whether VLMs can detect subtle changes in facial features beyond skin tone. While darker skin tone is an important feature of perceived Blackness (Stepanova and Strube, 2012; Wilkins et al., 2010), future research may benefit from manipulations that can better isolate skin tone to enable a more controlled assessment of its impact on group representations in AI models.
Another limitation is the lack of generalizability to real, non-computer-generated faces. Although GANFD images are generally perceived as authentic, with 13.15% of raters identifying the 40 faces in this study as artificial, it is unclear if the findings apply to real-world faces. The use of GANFD images was essential for isolating skin tone effects, which is challenging with real-world faces where darker skin often correlates with other varying features, introducing confounds. GANFD images provided the necessary control by keeping other facial features consistent, allowing us to focus solely on the impact of skin tone on group representations.
Finally, a limitation of this work is that the open-source models did not consistently generate stories in response to the writing prompt, instead producing detailed descriptions of the individuals in the images. While this deviation highlights a failure to follow instructions, it also underscores the robustness of our findings. The documented biases were not confined to the intended story format but were also evident in the more objective visual descriptions generated by the open-source models. This suggests that the observed biases in group representations are pervasive and not merely artifacts of specific text formats, further highlighting the broader implications of our results.
6 Conclusion
We find that Vision-Language Models (VLMs) reproduce homogeneity bias, particularly in relation to skin tone. By analyzing computer-generated images of Black American men and women with controlled skin tone variations, we found that VLMs produced more homogeneous stories for darker-skinned individuals in three of four models. Black women were consistently represented more homogeneously than Black men, with skin tone having a greater impact on women in two models. These results highlight how biases, including homogeneity bias, persist in multimodal AI systems, emphasizing the urgent need for research to address intersectional biases in AI.
Appendix S1 Appendix: GANFD Face Stimuli
Table S1: The identifiers of GANFD images used to represent the four groups.
Gender Skin Tone Image IDs
Men Lighter 1407-752417, 2308-489, 2491-1250, 3490-822, 4239-28, 7902-1229, 14792-3566, 19187-28, 22913-28, 24490-28
Men Darker 1407-3876, 2308-151, 2491-1407, 3490-3876, 4239-3876, 7902-38, 14792-533, 19187-533, 22913-533, 24490-533
Women Lighter 1402-28, 2617-2947, 10571-4022, 12360-28, 13372-1119, 13571-3112, 16252-2157, 17235-4022, 19933-533, 25885-28
Women Darker 1402-533, 2617-533, 10571-1407, 12360-533, 13372-533, 13571-3876, 16252-3876, 17235-3876, 19933-3876, 25885-115
Appendix S2 Appendix: Output of Mixed-Effect Models
Table S2: Summary output of the Skin Tone Models across all four VLMs. A significantly positive Skin Tone term indicates that cosine similarity was notably greater for darker-skinned individuals.
Skin Tone Model
GPT-4o mini GPT-4 Turbo Llama-3.2 BLIP-3
Fixed Effects
Intercept -0.0037 -0.053 -0.027 0.087
(0.041) (0.025) (0.036) (0.045)
Skin Tone 0.044∗∗∗ 0.15∗∗∗ $0∗∗$099∗∗∗
(0.0026) (0.0027) (0.0036) (0.0033)
Random Effects ($\mathbf{\sigma^{2}}$)
Pair ID Intercept 0.18 0.068 0.21 0.33
Residual 0.82 0.93 0.80 0.69
Observations 499,000 499,000 499,000 499,000
Log likelihood -657,971.70 -690,385.00 -651,912.20 -616,619.10
Table S3: Summary output of the Gender Models across all four VLMs. A significantly positive Gender term indicates that cosine similarity was notably greater for women than for men.
Gender Model
GPT-4o mini GPT-4 Turbo Llama-3.2 BLIP-3
Fixed Effects
Intercept -0.30 -0.053 -0.18 -0.084
(0.039) (0.025) (0.047) (0.063)
Gender 0.63∗∗∗ 0.15∗∗∗ 0.40∗∗∗ 0.24∗∗∗
(0.055) (0.0027) (0.067) (0.089)
Random Effects ($\mathbf{\sigma^{2}}$)
Pair ID Intercept 0.083 0.068 0.17 0.32
Residual 0.82 0.93 0.80 0.69
Observations 499,000 499,000 499,000 499,000
Log likelihood -658,069.90 -690,385.00 -652,143.80 -617,052.80
Table S4: Summary output of the Skin Tone and Gender Interaction Models across all four VLMs. A significantly positive interaction effect indicates that the effect of skin tone was significantly greater for women than for men.
Skin Tone and Gender Interaction Model
GPT-4o mini GPT-4 Turbo Llama-3.2 BLIP-3
Fixed Effects
Intercept -0.29 -0.17 -0.17 -0.036
(0.039) (0.032) (0.047) (0.063)
Skin Tone -0.0070 0.16∗∗∗ -0.023∗∗∗ -0.095∗∗∗
(0.0036) (0.0039) (0.0051) (0.0047)
Gender 0.58∗∗∗ 0.23∗∗∗ 0.29∗∗∗ 0.24∗∗
(0.055) (0.045) (0.067) (0.089)
Interactions 0.10∗∗∗ -0.0048 0.21∗∗∗ -0.0072
(0.0051) (0.0055) (0.0071) (0.0067)
Random Effects ($\mathbf{\sigma^{2}}$)
Pair ID Intercept 0.083 0.056 0.18 0.32
Residual 0.82 0.93 0.80 0.69
Observations 499,000 499,000 499,000 499,000
Log likelihood -657,739.80 -690,379.70 -651,482.60 -616,620.40
Table S5: Results of likelihood-ratio tests.
Model Term $\mathbf{\chi^{2}}$ p
Skin Tone 289.55∗∗∗ <.001
GPT-4o mini Gender 87.52∗∗∗ <.001
Interaction 389.86∗∗∗ <.001
Skin Tone 3206.77∗∗∗ <.001
GPT-4 Turbo gender 22.86∗∗∗ <.001
Interaction 0.76 .38
Skin Tone 502.39∗∗∗ <.001
Llama-3.2 Gender 32.36∗∗∗ <.001
Interaction 838.42∗∗∗ <.001
Skin Tone 881.34∗∗∗ <.001
BLIP-3 Gender 7.34∗∗ .007
Interaction 1.17 .28
Table S6: The effect of skin tone within each gender group across all VLMs.
Model Gender Effect of Skin Tone 95% CI
GPT-4o mini Men -0.0038 [-0.058, 0.050]
Women 0.040 [-0.014, 0.094]
GPT-4 Turbo Men -0.053 [-0.097, -0.0085]
Women 0.10 [0.058, 0.15]
Llama-3.2 Men -0.027 [-0.092, 0.038]
Women 0.053 [-0.012, 0.12]
BLIP-3 Men 0.087 [-0.00051, 0.17]
Darker -0.012 [-0.10, 0.075]